OpenStudy (anonymous):

Shannon makes large cookies. These cookies are 22 inches in diameter. How much icing will she need to cover the top of each cookie? 11π in2 22π in2 484π in2 121π in2

OpenStudy (bahrom7893):

Find the area of each cookie: d=22 =>r=11 A=pi*r^2=pi*11^2=121pi

OpenStudy (anonymous):

Gotcha , so 121